OpenStudy (anonymous):

what's the integral of (1+e^x)/(1-e^x) ? I come up with the answer, x + 2ln|1-e^x| + C I'm not sure if my answer is correct. Somebody please help?

OpenStudy (anonymous):

\[{1+e^x \over 1-e^x}={1-e^x+2e^x \over 1-e^x}=1+2{e^x \over 1-e^x}.\] The integration of is \(x\), and the integration of \(2\frac{e^x}{1-e^x}\) is \(-2\ln|1-e^x|\), by substitution.

OpenStudy (anonymous):

of 1 is \(x\)*

OpenStudy (anonymous):

Then you add +constant.

OpenStudy (anonymous):

okay so my answer was correct. thanks!